

********************************************************************
UARS WEEKLY STATUS REPORT - 07 Jan 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 63% of
each day's data amounting to nearly 556 minutes of solar observing time
per day.
***********
SUSIM
***********
During this past week, the SUSIM experiment successfully gathered
daily mid (1.1 nm) resolution, MgII, occultation, solar pointing
offset, and continuous monitoring scans for time periods with UARS
telemetry or tape recording.
***********
PEM
***********
The new PEM2K:: is now fully configured and functional. All communication
paths should function as normal.
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004239 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004240 and 2004249 are currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC archive
are between 1991274 and 2004229. MEPS data between 2004230 and 2004239
are waiting for archival at the DAAC.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 25 December 2004 (Orbit 72729, GMT Day 360) through
31 December 2004 (Orbit 72834, GMT Day 366).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 361/0032 72744 QB/PB B ON
PEM 361/2249 72758 PEM AXIS 1 & 2 OFF
PEM 361/2347 72759 ZEPS/NEPS DAY-ONLY
PEM 363/1401 72783 PEM TO FULLTIME
PEM 363/1401 72783 PEM AXIS 1 & 2 ON
PEM 365/2236 72818 Z/H FULLTIME TO DAY-ONLY
The beta angle ranged between the angles of 76.7 degrees to 52.3
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 2.0 degrees on 12 January 2005 (DOY 012).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.64 degrees C. Additional
battery performance data for each battery on 24 December and 31
December 2004 are provided in Appendix D.
The clock error ranged from 07.7 to -11.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 05
WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued their CMS software acceptance testing this week.
This testing will continue into the new year. The goal for completion
is no later than end of January 2005.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 January 2005 (DOY 012).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 31 December 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
72744 361/0032 04-360 WINDII QB/PB B ON (S/C POWER MGT)
72753 361/1445 04-361 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
72758 361/2249 04-362 PEM PEM AXIS 1 & 2 OFF
72759 361/2347 04-363 PEM ZEPS/NEPS DAY-ONLY
72773 362/2202 04-364 MPS CHANGE V/T LIMIT FROM LV5 TO LV6
72783 363/1401 04-366 PEM PEM TO FULLTIME
72783 363/1401 04-365 PEM PEM AXIS 1 & 2 ON
72818 365/2236 04-367 PEM PEM (Z/H) FROM FULLTIME TO DAY ONLY
72820 366/0052 04-368 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 31 December 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 31 December 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
72753 361/1510 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
72820 366/0057 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 31 December 2004
24 December 2004 - GMT Day 359 Beta = 79.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.48/-2.42 +00.0/+00.0 07.0 0.00 0.0
2 -11.2/ -44.8 +2.28/+1.65 +11.6/-12.8 27.7 1.08 9.6
3 -72.8/-184.8 +1.65/+0.71 +11.2/-12.0 27.7 1.07 9.2
31 December 2004 - GMT Day 366 Beta = 79.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.62 +00.0/+00.0 07.0 0.00 0.0
2 44.8/ -11.2 +4.51/+3.87 +23.2/-14.4 26.7 1.05 16.7
3 -11.2/ -89.6 +3.23/+2.60 +22.0/-13.6 26.7 1.04 16.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
72773 362/2202 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 31 December 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-090 FOT 72780 363/0906 30832 Generic Late Acquisition
04-090 FOT 72793 364/0612 30832 Generic Late Acquisition
04-094 FOT 72814 365/1550 30853 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-091 FOT 72530 346/1854 37847 No UPD/GCMR Capability
04-092 FOT 72635 353/1831 37983 Late Acq. - TDRS West
04-093 FOT 72717 359/0407 37991 No Acq. - TDRS 275
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-090 FOT See List Dec 04 30832 Generic Late Acquisition
04-094 FOT 72814 365/1550 30853 Late Acquisition